OpenStudy (anonymous):

6690 9900 what are three common factors?

OpenStudy (karatechopper):

2,5,10

OpenStudy (anonymous):

what about prime factors?

OpenStudy (karatechopper):

2,3, 5

OpenStudy (anonymous):

\[6690=2^1*3^1*5^1*223^1 \]\[9900=2^2*3^2*5^2*11^1 \]Karatechopper is correct